| |
Minutes of the Mechanical Group Meeting
|
Date |
9 November 2000 |
| |
|
Attendance |
K Dee, P Morrall, M V D Hoeven, M Blanken, P
Jolley, S Rodreigez (part time) |
| |
|
No |
Details |
Action |
|
1 |
Minutes of the previous meeting
4/10/00 |
|
| |
These were accepted as a true record |
|
| |
|
|
|
2 |
Action items |
|
|
2.1 |
NAOMI Trolley |
|
| |
Requires a cover made over the whole table
speadyframe ordered. Check prices for Perspex |
PSM |
| |
Requires breaks - no progress |
SBR |
| |
Requires a dedicated lifting frame - on
order |
PSM |
| |
Following the recent instrument change, the
procedures need updating - no progress |
JOLLEY |
| |
Removal of INGRID from NAOMI is difficult. New
lifting arrangement for INGRID to be defined - no progress |
KMD |
| |
|
|
|
2.2 |
Cable Wrap |
|
| |
Document containing cable descriptions and
positions to be updated. CMWJ to be contacted - no progress |
KMD |
| |
|
|
|
2.3 |
Second Optical GHRIL Table |
|
| |
Following recent problems with the old optical
table with SCAM, the provision of a more stable table to be
reviewed. New table to be ordered. |
KMD |
| |
|
|
|
2.4 |
INGRID |
|
| |
Instrument change schedules to reflect the
following:
Moving INGRID from cass to GHRIL 3 days
Moving INGRID from GRHIL to cass 4 days - complete |
KMD |
| |
Duty checks to include a test for sweating - complete |
JOLLEY |
| |
INGRID bracket needs reviewing to include
easier initial alignment and fine adjustments |
KMD |
| |
INGRID change document to be updated - complete |
MFB |
| |
|
|
|
2.5 |
GRHIL |
|
| |
GRHIL roof safe working to be reviewed.
Harnesses to be used - paper note to be distributed |
KMD |
| |
GRHIL roof to have dedicated strops - to be
ordered |
PSM |
| |
|
|
|
2.6 |
Snow cleaning |
|
| |
Need a regime for snow cleaning. MFB to ensure
its scheduled and completed with the aid of the telescope manager
and ops team leader. |
MFB |
| |
Note to be issued clarifying the importance of
snow cleaning and the new procedures for doing so. |
MFB/ KMD |
| |
|
|
|
3 |
ING Overview |
|
| |
KMD reported that the during the recent ING
Boards visit, they were positive of the gains made in the previous
year. The ING should be doing more development work with a view to
offering its services to external organisations. The board also
hinted that PPARC are still considering the prospect of joining
ESO. The NWO also see a future with the ING and intend to continue
there support over the coming years. |
|
| |
|
|
|
4 |
Project Reports |
|
|
4.1 |
INT Counterbalance |
|
| |
Requires some modifications to allow working in
all orientations. New gearboxes to be fitted and protective covers
required. |
PSM |
| |
PSM to explain instrument change procedure to
SRB. |
PSM |
| |
PSM to replace centre screws |
PSM |
| |
PSM to install protective rubber bounce pads |
PSM |
| |
|
|
|
4.2 |
Safety |
|
| |
Jolley to convert and publish test certificates
on the intranet |
JOLLEY |
| |
PSM to liase with AKC to update the lifting
document on the intranet. The new document to include authorised
lifting personnel, the equipment that they can use and there
training requirements/schedules. |
PSM |
| |
Lifting Store needs sorting out. PSM and JC to
implement a routine for everyone to follow |
PSM/JC |
| |
|
|
|
4.4 |
Red CCD's |
|
| |
David's manufacturing these items expected
complete mid December |
David |
| |
|
|
|
4.5 |
Small Fibres |
|
| |
Prism alignment going well. Jig to be finished.
Test to be performed |
MVDH/ MFB |
| |
New fingers/slit unit to be ordered. |
JOLLEY |
| |
MFB/JOLLEY to arrange coating of the
microlenses |
MFB/ JOLLEY |
| |
Transmission tests to be finalised |
MFB |
| |
|
|
|
4.6 |
OASIS |
|
| |
This is a new project to the group, so JOLLEY
gave a brief explanation of what it is and what we need to
provide. |
|
| |
|
|
|
4.7 |
Other Projects |
|
| |
PN Spectrograph - due next semester |
|
| |
LIRIS - due soon |
|
| |
Auxport Camera - no resources available |
|
| |
Long Camera - waiting for optical design and
resources |
|
| |
Near Earth Object (NEO) - close to getting
funding for JKT |
|
| |
|
|
|
5 |
Maintenance |
|
|
5.1 |
Oil system at the WHT needs inspecting and
possible cleaning |
SRB/ KMD |
| |
|
|
|
5.2 |
For information only: KMD stated that care
should be taken when working around the WHT bull ring. The
microswitch has three positions and it is possible to drive the
telescope passed its end stops. IF YOU THINK THIS HAS BEEN
KNOCKED/MOVED THEN CHECK BEFORE RELEASING TELESCOPE. |
|
| |
|
|
|
5.3 |
Following the recent request the WHT secondary
mirror alignment was checked and found to be acceptable. |
|
| |
|
|
|
6 |
Training |
|
| |
No reports |
|
| |
|
|
|
7 |
Administration |
|
| |
No Report |
|
| |
|
|
|
8 |
A.O.B |
|
|
8.1 |
Mechanical Design Support |
|
| |
Kevin to source extra short term design
support. |
KMD |
| |
|
|
|
8.2 |
Requests for any work that is deemed to be an
enhancement must come from the EMM and not individuals.
Work that is not an enhancement but necessary for the working of
the telescope to be carried out as normal. |
ALL |
| |
|
|
|
9 |
Next Meeting |
|
| |
TBD |
|